### v1

# Understanding How Psychological Distance Influences User Preferences in Conversational Versus Web Search

Abstract.
Conversational search offers an easier and faster alternative to conventional web search, while having downsides like lack of source verification. Research has examined performance disparities between these two systems in different settings. However, little work has considered the effects of variations within a given search task. We hypothesize that psychological distance - one’s perceived closeness to a target event - affects information needs in search tasks, and investigate the corresponding effects on user preferences between web and conversational search systems. We find that with greater psychological distances, users perceive conversational search as more credible, useful, enjoyable, and easy to use, and demonstrate increased preference for this system. We reveal qualitative reasons for these differences and provide design implications for search system designers.
Conversational Search, Web Search, Information Seeking, Information Retrieval, User Experience, Psychological Distance, Construal Level Theory, Large Language Models, Generative AI
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Information systems Search interfacesccs: Computing methodologies Artificial intelligenceccs: Information systems Web searching and information discovery
1. Introduction
As digital technologies continue to advance, the ways people interact with information have significantly evolved. Users are no longer confined to traditional web search engines like Google Search (Kobayashi and Takeda, 2000; Brin and Page, 2012), but increasingly have the option of conversational search interfaces powered by Large Language Models (LLMs) like ChatGPT (Ai et al., 2023; Liao et al., 2020). Conversational search systems can potentially increase usability (Tariverdiyeva, 2019), enhance efficiency (Ai et al., 2023) by synthesising information from extensive underlying datasets (Deng and Lin, 2022), and increase user enjoyment through humanlike dialogues (Rhim et al., 2022; Følstad and Brandtzaeg, 2020; Brandtzaeg and Følstad, 2017). However, such systems also have certain downsides: hallucination of facts and information (Ji et al., 2023) and lack of source attribution can reduce perceptions of information completeness and verifiability (Jung et al., 2024), negatively affecting user trust, acceptance, and system adoption. Prior work has investigated user preferences for these two information retrieval (IR) systems in a range of tasks. Conversational search was found to provide superior information to web search in informal medical diagnosis (Sandmann et al., 2024; Van Bulck and Moons, 2024; Shen et al., 2024), and lower cognitive load (Kumar et al., 2024) and greater efficacy (Arias Sosa and Godow, 2023) in coding-related queries.
However, a crucial overlooked factor is the inherent variation in individual tasks. Notably, tasks can vary significantly in terms of the perceived psychological distance to the user: for instance, whether they are seeking information for themselves or someone else, or whether they are searching about an event more or less likely to happen. The well-established Construal Level Theory (CLT) describes the relation between psychological distance and the level of abstraction of thought (Trope and Liberman, 2010). The psychological distance of a task (e.g. planning a trip for next week versus next year) influences whether people think about specific details (e.g. public transport options) or general ideas (e.g. cities to visit) (Trope and Liberman, 2010). Such variations significantly impact how users process information and make decisions (Lermer et al., 2015; Trope and Liberman, 2010; Wakslak and Trope, 2009; Fiedler, 2007); we propose that this effect extends to information needs in IR tasks, and thus user perceptions and preferences of the IR systems that fulfil them. Psychological distance is a common factor in everyday life, being assessed automatically without conscious effort (Liberman and Trope, 2014), and has fundamental and widespread importance for cognitive and behavioral decision making (Liberman and Trope, 2014; Maglio et al., 2013). Therefore, the effect of psychological distance of IR tasks on preferences for IR systems presents a key research gap.
To investigate this, we conducted a 2x4 between-subjects study, with 128 participants divided into eight groups. Each participant was assigned to one of four psychological distance dimensions posited by CLT — spatial (where an event takes place), temporal (when an event occurs), social (who is involved in the event), or hypothetical (how likely the event is to happen) — and one of two psychological distance levels (near or far). Participants completed an IR task, manipulated according to their assigned dimension and distance, using both web and conversational search. They were then asked to rate the perceived credibility, perceived usefulness, enjoyability, and ease of use of each tool, and finally their preference between tools.
Our findings indicated that participants in the far psychological distance condition exhibited a significantly higher level of preference for the LLM tool over traditional web search, reporting greater perceived credibility, perceived usefulness, enjoyment, and ease of use. Additionally, participants’ search behaviors varied depending on psychological distance. When distance was perceived as greater, participants more effectively utilized the strengths of the LLM system to explore broad ideas, such as cities, cuisines, and accommodations. In contrast, nearer distances prompted participants to focus on concrete steps for planning and executing their trips. These findings suggest that the choice of IR system is closely tied to the user’s situational context, and offer important design implications for designers of IR systems.
Our work makes the following contributions:
  1. (1)
    It advances our understanding of the differences in user perceptions and preferences when interacting with information through traditional search engines and LLM-powered conversational search, contributing to the literature on Human-Information Interaction. Additionally, we explore how these preferences can inform tool recommendations based on psychological distance for information-seeking.
  2. (2)
    It offers valuable insights for optimizing future IR systems to better align their functional designs with varying user needs across diverse contexts, ultimately enhancing user experience and satisfaction in an increasingly complex digital landscape. Notably, our findings suggest that conversational search systems should be prioritized for tasks involving greater psychological distance, as this factor significantly influences user acceptance of IR technologies.
2. Related Work
2.1. Information Seeking
Information seeking originates from the library sciences, but has been increasingly studied in the context of computer science (Blandford and Attfield, 2022). Information seeking activities comprise three main processes: experiencing information needs, information seeking, and information use (Blandford and Attfield, 2022; Choo et al., 2013). The former two are relevant to our current study. Information needs arise from problems encountered in specific situations and are influenced by multiple factors, such as individual knowledge gaps (Allen, 1996), the level of information need (Taylor, 1968), and the constraints of time and resources (Choo et al., 2000). Information seeking involves ”identifying, selecting, and interacting with information sources to purposely retrieve information” (Choo et al., 2013). Each specific information seeking task influences both information needs, which in turn affects an individual’s choice of information sources while looking for information (Choo et al., 2000), and information seeking, where the nature of the task affects the amount of effort and psychological state invested in the task (Cole, 2011; Byström and Järvelin, 1995). Users gravitate toward information sources that are most relevant to their needs (Case and Given, 2016), and the personal sense of control that comes from finding information that matches their needs encourages further use of its source (Wilson, 1997). In sum, the nature of an information seeking task affects a user’s choice of information source, and hence their choice of information seeking tool. This motivates our study of how the nature of tasks affects preferences for different IR systems.
2.2. Construal Level Theory and Psychological Distance
To explore how specific tasks influence user preferences for web or conversational search, we turn to CLT, which provides a theoretical framework for manipulating task situation through psychological distance. Psychological distance is a common and critical factor in information seeking tasks, such as planning for near or distant events, searching information for oneself or someone else. These variations in psychological distance can significantly impact how users process information and make decisions (Lermer et al., 2015; Trope and Liberman, 2010; Wakslak and Trope, 2009; Fiedler, 2007). CLT not only provides ways to manipulate psychological distance of tasks, but also provides a theoretical basis for the potential impact on user preferences towards different IR systems. When users make choices or set preferences for different objects, they rely on their understanding of these objects (Trope and Liberman, 2010). In CLT, increasing psychological distance enhances higher-level understanding, which in turn influences user evaluations and preferences (Trope and Liberman, 2010). CLT posits that as the psychological distance of a mental phenomenon increases, the construal level of thought processes shifts from low-level concrete constructions to high-level abstract constructions (Trope and Liberman, 2010). We believe that in the context of information seeking, the perceived psychological distance of an information seeking task presented to a user affects their construal level of thinking. Task psychological distances may be smaller (e.g. finding information for oneself) or greater (e.g. finding information for someone else). CLT also discusses how psychological distance affects task-related evaluations (Trope and Liberman, 2010). As psychological distance increases, users tend to seek more supportive viewpoints than opposing ones (Trope and Liberman, 2010). This is because higher-level thinking prioritizes the pros of a situation over its cons (Trope and Liberman, 2010). Herzog et al. (Herzog et al., 2007) found that as psychological distance increases, it becomes easier for users to generate supportive viewpoints and harder to generate opposing ones. Thus, we propose that by varying the psychological distances of tasks, users will perceive the nature of their task differently via different construal levels(Trope and Liberman, 2012), and hence tend toward either web search or conversational search.
CLT proposes four main dimensions of psychological distance: spatial, temporal, social, and hypothetical (Trope and Liberman, 2010). Spatial distance is where events are going to happen; temporal distance concerns when events are going to happen; social distance is how similar the people involved in an event are to oneself; and hypothetical distance is the likelihood of an event happening (Trope and Liberman, 2010). Each dimension has some unique characteristics, but all four influence the mental models of a user through their construal level (Liberman and Trope, 2014). Low-level constructions are unstructured and detail-rich, encompassing primary, secondary, and incidental characteristics of an event (Trope and Liberman, 2010). In contrast, high-level constructions are more abstract and less specific, extracting essential points from available information while omitting details (Trope and Liberman, 2010). In comparing web search and conversational search, we propose that web search aligns more with the need for low-level constructions, while LLM-powered conversational search meets higher-level construction needs. Therefore, our study is based on the theory that the psychological distance associated with information seeking tasks influences user information need through different levels of construal. The information seeking activities users engage in to satisfy their current information needs, in turn, affect their perceptions and preferences for these two technologies.
2.3. Conversational Search as a New Means of IR
IR is the task of searching for relevant information from an online repository to fulfil a user need. Conversational search, a specialized branch of IR systems, aims to help users obtain information that meets their needs through flexible and natural dialogues, based on their queries and relevant feedback (Kiesel et al., 2021; Jannach et al., 2021; Liao et al., 2020; Radlinski and Craswell, 2017; Croft and Thompson, 1987). Modern LLMs provide a revolutionary step forward in empowering such systems and their advantages, due to their outstanding performance in understanding and generating human-like text (Gao et al., 2020).
Conversational search offers a number of innate advantages over traditional web search. Unlike web search queries that require specific terms and keywords, conversational search enables users to express their needs through natural dialogue, reducing the effort needed to clarify their particular aims (Radlinski and Craswell, 2017). Conversational systems also significantly reduce cognitive load (Kaushik, 2019) by providing summaries instead of the raw content of traditional search engines (Kaushik and Jones, 2023). For search tasks that involve multiple query aspects, such as making travel plans that include accommodation, food, and attractions, conversational search can be preferable in allowing users to assess the overall utility of such multi-item compositions (Radlinski and Craswell, 2017). Furthermore, the natural conversational abilities and human-like qualities of LLMs provide greater social interaction, which itself is valuable in enhancing the overall search experience (Følstad and Brandtzaeg, 2020).
However, LLM-powered conversational search also suffers from a few disadvantages. LLMs, as natural language generation models (Brown et al., 2020), inevitably experience hallucination (Maynez et al., 2020), generating information that contradicts or is not supported by their source content (Ji et al., 2023). Moreover, LLMs often do not provide specific sources for the summaries they generate (Bender et al., 2021). Finally, because their training data are based on past internet information with a specific cut-off point, (Achiam et al., 2023), the information provided by LLMs can be outdated, and they may be unable to answer queries that require the latest data or information (Liu et al., 2023). This therefore raises the question of when LLM may or may not be superior to traditional web search.
2.4. Psychological Distance in Information Seeking Tasks
Previous research has highlighted the potential advantages of conversational search in terms of ease of use (reducing user effort), usefulness (especially for complex tasks), and enjoyability, as well as potential issues with credibility. However, our exploration of user preferences for conversational versus web search considers not only the impact of the interactive aspect, but also the nature of the information seeking tasks. As mentioned in section 2.1, the nature of information seeking tasks influence users’ choices of information sources, which should affect their preferences for different IR systems.
Previous studies have investigated the relative efficacy of conversational search and web search across various types of information seeking tasks, such as healthcare and tourism. Changes in task type can alter user information needs, with conversational search and web search differing in their ability to meet these needs. Specifically, in the healthcare context, studies have demonstrated that generative pre-trained transformers (GPTs) surpass traditional web searches like Google in diagnosing prevalent illnesses and suggesting subsequent treatment steps (Hristidis et al., 2023; Shahsavar et al., 2023; Sandmann et al., 2024). In the realm of tourism, ChatGPT has been shown to significantly enhance the planning of trips. They serve as virtual assistants, offering personalized advice, and contribute to a richer travel experience by integrating diverse information sources (Dwivedi et al., 2024; Zarezadeh et al., 2023; Wong et al., 2023; Zhang et al., 2024). Despite these advancements, these studies also highlight the limitations in conversational search, closely related to the challenges associated with LLMs mentioned in section 2.3, underscoring the importance of traditional web search (Sandmann et al., 2024). Users frequently turn to web search to verify LLM-generated information, indicating that LLMs complement rather than replace web search in information seeking tasks (Yen et al., 2024).
However, even if the type of task remains constant, psychological distance, a common factor in information seeking tasks, can also lead to changes in user information needs. Few studies have explored the impact of psychological distance in information seeking tasks. Moreover, previous studies have rarely focused on user perceptions and preferences for these two IR systems, lacking a comprehensive perspective from task nature to technology preference. Therefore, this study aims to fill this gap by considering the psychological distance in information seeking tasks to deepen our understanding of user preferences for conversational search versus web search in real-world information searching. This not only helps us to understand the motivations behind user choices at specific task psychological distances, but also provides important insights for designing IR systems that better meet user needs. Since CLT provides a theoretical foundation for both manipulating tasks through the critical factor (psychological distances) and understanding their impact on user preferences, we investigate this gap by manipulating the psychological distance of information seeking tasks using CLT. We are particularly interested in exploring:
  • RQ: How psychological distance of information seeking tasks influences user preferences for web search or LLM-powered conversational search.
Web search conventionally provides a list of results ranked according to the relevance of the user’s query, allowing users to access specific information through links (Brin and Page, 1998). However, obtaining a comprehensive overview of a topic is more difficult as users must access and filter through multiple sources and aggregate this information themselves. Conversely, LLM-powered conversational search generates summaries of the information relevant to the user’s current search intent, providing a more comprehensive overview of relevant information. However, these tend to lack specific details (Achiam et al., 2023; Brown et al., 2020; Devlin et al., 2018), and require user to ask further questions or refine their prompts for more detailed inquiries. Further, when users require highly-specific or time-sensitive information, LLM-powered conversational search may not provide accurate answers (e.g. local train ticket prices or current weather conditions).
CLT posits that tasks with greater psychological distances elicit preferences for more abstract, higher-levels overviews of information. Thus, we propose that the perceived ease of use of each IR system depends on the psychological distance of the task and thus the needs of the user. We hypothesize:
  • H1: Users engaging in information seeking tasks with greater psychological distances perceive greater ease of use with LLM-powered conversational search compared to web search, relative to users engaging in tasks with smaller psychological distances.
CLT suggests that as psychological distance increases, users tend to seek more supportive rather than opposing viewpoints, due to high-level thinking’s tendency to overlook specific disadvantages. In conversational search, recent studies (Sharma et al., 2024) suggest that LLM-powered conversational search creates an echo chamber effect, producing content that validates user views more than web search and reducing the occurences of opposing viewpoints. Content that corroborates user views is perceived as more credible (Nickerson, 1998), specifically in terms of accuracy, authenticity, and believability (Appelman and Sundar, 2016). Therefore, we hypothesize the following.
  • H2: Users engaging in information seeking tasks with greater psychological distances perceive higher informational credibility in LLM-powered conversational search compared to web search, relative to users engaging in tasks with smaller psychological distances.
Technology Acceptance Models (TAMs) (Davis, 1989; Davis et al., 1992; Ayeh, 2015) offer a theoretical basis for exploring the user acceptance and other related perceptions of conversational search by examining user perceptions of the ease of use and information credibility. TAM posits that two primary factors influence a user’s decision to adopt a technology: perceived usefulness and perceived ease of use (Davis, 1989). These perceptions directly influence users’ attitudes towards the technology, while ease of use also indirectly affects attitudes by influencing perceived usefulness (Adams et al., 1992).
Figure 1. Extended TAMs incorporating perceived enjoyment and information credibility. This model illustrates the relationships between perceived ease of use, perceived usefulness, perceived enjoyment, and information credibility with the acceptance intention of technology.
[…]
This diagram visually represents an extended Technology Acceptance Model (TAM). It consists of four main rectangular components, each labeled with a different factor influencing technology acceptance. These components are connected by arrows indicating the direction of influence, each marked with a plus sign to denote positive relationships. Starting from the left, the first component is labeled ”Perceived Information Credibility”, which connects directly to ”Perceived Usefulness”, a central component in the middle of the diagram. Above ”Perceived Usefulness” is ”Perceived Ease of Use”, which influences ”Perceived Usefulness”. Below ”Perceived Usefulness” is ”Perceived Enjoyment”, linked to it by a bidirectional arrow, indicating a mutual positive influence between the two. ”Perceived Ease of Use”, ”Perceived Usefulness” and ”Perceived Enjoyment” all have direct connections to ”Acceptance Intention”, depicted as an oval on the far right, indicating the final outcome of the model. The diagram uses a color scheme where ”Perceived Usefulness” and ”Perceived Enjoyment” are highlighted in light green, while ”Perceived Information Credibility” and ”Perceived Ease of Use” are in light blue, representing a special focus on ”Perceived Information Credibility” and ”Perceived Ease of Use”.
Extensions of the TAM have identified additional relevant variables beyond perceived usefulness and ease of use. One such factor is perceived enjoyment (Davis et al., 1992) expanded the TAM to include perceived enjoyment, which can also directly enhance user attitudes and perceived usefulness: when users find a technology enjoyable, they also perceive it as more useful, and vice versa. In another extension of the TAM for consumer-generated media that aids travel planning, Ayeh (Ayeh, 2015) introduced information credibility as a prerequisite for perceived usefulness. Applying the TAM and its extensions in conversational search, we illustrate the interactions among user perceptions of these systems in Figure 1. Based on H1 and H2, we propose the following hypotheses:
  • H3: Users engaging in information seeking tasks with greater psychological distances perceive higher usefulness in LLM-powered conversational search compared to web search, relative to users engaging in tasks with smaller psychological distances.
  • H4: Users engaging in information seeking tasks with greater psychological distances perceive higher enjoyment in LLM-powered conversational search compared to web search, relative to users engaging in tasks with smaller psychological distances.
  • H5: Users engaging in information seeking tasks with greater psychological distances have higher acceptance in LLM-powered conversational search compared to web search, relative to users engaging in tasks with smaller psychological distances.
3. Methodology
Our experiment comprised eight groups, involving the four dimensions of psychological distance (spatial, temporal, social, and hypothetical) each with two levels of distance (near and far), constituting a 4x2 design. Participants in each experimental group used two IR technologies: a web search (WebSearch) system and an LLM-powered conversational search (ConvSearch) system.
3.1. Task Design
In the experiments, participants were asked to take an information seeking task with fictional travel scenarios. In the tasks, participants were asked to imagine planning a one-week vacation to another country, focusing on food, natural attractions, and historical buildings.
We varied the psychological distance between the participant and the task as the main independent variable (two levels: near or far) in a between-subjects design, and repeated this across four dimensions of psychological distance (spatial, temporal, social, or hypothetical). These designs allowed us to explore whether there are differences in user perceptions and preferences towards IR technologies based on psychological distance, and whether these differences are consistent across the different dimensions or influenced by other factors.
Table 1. Manipulations of Psychological Distance Across Four Dimensions in Information Seeking Tasks
Dimension Spatial Temporal Social Hypothetical
Manipulation Trip destination Trip timing Traveller identity Likelihood of trip occurring
Psych. Dist. Task Description
Near Nearby country Next week Close friend of participant Confirmed, plane tickets booked
Far Faraway country Next year Internet user on online forum Not confirmed, only a possibility
Neutral Nearby country Next month The participant themselves (Not mentioned)
We varied a particular aspect of the tasks  [^1] Details on the specific scenarios are fully outlined in Appendix 7.1. according to the psychological dimension being manipulated in each experimental group, as seen in Table 1. We introduced a ”neutral” setting in our study, which does not constitute an independent experimental condition. Instead, it ensures that dimensions of psychological distance other than the one currently under focus remain neutral. This prevents these dimensions from influencing participants’ perception of the psychological distance being studied. For example, in ”neutral” distances, the choice of a nearby country of the residence of our participants as a destination reflects the reality that most international travel typically involves neighboring countries (McKercher and Mak, 2019; McKercher et al., 2008). Similarly, in describing the traveler in dimensions other than social distance, we choose the participant themselves because conducting information searches and planning trips for oneself is more common in everyday life.
Each participant was required to use both a WebSearch and a ConvSearch to perform the same tasks for different destinations, so that the first search using WebSearch did not influence the subsequent search with ConvSearch. In the same experimental group, the difference in destination did not affect psychological distance: Nearby countries in the set of near or neutral spatial distances were represented by Mexico (Information Seeking Task 1) and Cuba (Information Seeking Task 2), both geographically close to our participants in the United States, while faraway countries in the set of far spatial distances were represented by Egypt (Information Seeking Task 1) and Sudan (Information Seeking Task 2). These countries were chosen to minimize the influence of prior knowledge by excluding mainstream American travel destinations and focusing on non-Western countries.
3.2. Procedure
Figure 2. Flowchart of the study procedure detailing the sequence of tasks and surveys conducted across two information seeking tasks. This diagram illustrates the comprehensive steps from the initial pre-task survey through two distinct information seeking tasks, followed by post-task surveys and a final preference survey.
[…]
This figure presents a detailed flowchart of the study procedure, organized into six main phases across a horizontal layout. Each phase is represented as a rectangular box containing lists of activities or survey components, with arrows indicating the sequence from left to right. Step 1 is Pre Task Survey: This initial phase assesses participants’ baseline attitudes AI. Step 2 is Information Seeking Task 1: This phase starts with reading a scenario, followed by using the WebSearch to gather information. Participants’ search behaviors are recorded, and they are required to complete a travel plan. Step 3 is Post Task 1 Survey: This survey evaluates participants’ perceptions after the first task, focusing on Perceived Ease of Use, Usefulness, Enjoyment, and Information Credibility. Step 4 is Information Seeking Task 2: Similar to the first task, participants read a new scenario and use a Conversational Search. This phase also involves recording the conversation and completing a travel plan. Step 5 is Post Task 2 Survey: This survey mirrors the first post-task survey but evaluates the conversational search experience, assessing the same four perception metrics. Preference Survey: The final phase involves a survey where participants choose their preferred information retrieval system between WebSearch and the conversational search. They also provide reasons for their choice.
The full study design (Figure 2) includes six phases. In Step 1, participants completed a pre-task survey to measure their baseline attitudes to artificial intelligence (AI). In Step 2, the scenario for the first information seeking task was introduced. Participants then used WebSearch to seek information. We placed WebSearch first as it is a widely recognized IR system, and serves as a baseline for the experiment. To monitor the search behaviors of participants, we created a custom web search interface that integrated Google’s search API  [^2] https://developers.google.com/custom-search/v1/overview. After gathering the necessary information for the task, participants were required to write a travel plan consisting of at least 100 words to ensure the task completion quality. Step 3 was the first post-task survey, where participants were asked to rate their perceptions of ease of use, usefulness, enjoyment, and information credibility when using WebSearch. Steps 4 and 5 mirrored the previous two steps. Participants used ConvSearch to perform the second information seeking task, which had a different destination that had the same spatial distance as the first one. We built the ConvSearch system on the UChat [^3] https://uchat.au/ platform using OpenAI’s GPT-4 API  [^4] https://platform.openai.com/docs/api-reference/completions/object, similar to ChatGPT. Participants used our tool to seek information and then reported their perceptions of this technology in the second post-task survey. Finally, in Step 6, participants indicated their preference between the two IR systems they used and elaborated on reasons for their choice.
3.3. Participants
Participants were recruited on the CloudResearch Connect platform [^5] https://www.cloudresearch.com/products/connect-for-researchers/, with residency restricted to the United States to ensure uniformity in their perception of spatial distances to designated travel destinations. After excluding individuals who failed attention checks, the study consisted of 128 participants according to the power analysis result (16 per group; 62 females, 64 males, 2 undisclosed). The average age of the participants was 37.6 (SD = 10.5), and 63% of the participants had at least a bachelor’s degree. The study was expected to last 30 minutes, with participants compensated at CloudResearch’s recommended rate of \$10 per hour. This research was reviewed and approved by the Anonymous University Departmental Ethics Review Committee.
3.4. Measures
3.4.1. User perception, behavior and preferences
The main dependent variables were based on the TAM (Davis, 1989) and its variants (Davis et al., 1992; Featherman and Pavlou, 2003; Lee, 2009). We aimed to understand how psychological distance influences user behavior and preferences for WebSearch and ConvSearch through affecting these user perception variables. Most perception variables involved several questions, each measured on a 7-point Likert scale ranging from ”Strongly disagree” (1) to ”Strongly agree” (7). The mean score reported by participants for all questions within each variable was used to represent the score of that variable.
  • Perceived ease of use and Perceived usefulness are gauged using questions from TAM (Davis, 1989). For ease of use, questions include, ”Learning to operate web search/LLM-powered conversational search would be easy for me” and ”I would find it easy to get web search/LLM-powered conversational search to do what I want to do” among a total of six questions. Similarly, perceived usefulness is measured with six questions, such as ”Using web search/LLM-powered conversational search in my job would enable me to accomplish my tasks more quickly” and ”Using web search/LLM-powered conversational search would improve my job performance”.
  • Perceived enjoyment is measured using three specific questions from a TAM variant (Davis et al., 1992), such as ”I find using web search/LLM-powered conversational search enjoyable” and ”The actual process of using web search/LLM-powered conversational search is pleasant”.
  • Perceived information credibility is represented through three specific dimensions: accuracy, authenticity, and believability (following (Appelman and Sundar, 2016)). We provide participants with three sliders, each ranging from 1 to 100, to indicate their perceptions of information credibility across three dimensions. We then calculate the mean of these three values to represent the overall perceived credibility of the message.
  • User preferences are determined by asking participants which tool they would choose for similar tasks in the future and by requiring them to explain their reasons for this choice.
3.4.2. Thematic Analysis About User Preferences Reasons
We conducted a thematic analysis (Braun and Clarke, 2006) of the reasons participants (n = 128) reported their tool preferences to provide user-centered design recommendations for IR tools. Initially, two researchers independently performed semantic (what the participants said) and latent (researchers’ interpretations of the data) analyses on the data to generate preliminary codes (without predefined categories). The research team refined these initial codes through multiple discussions and developed themes using a bottom-up inductive approach. Subsequently, two researchers coded the reasons participants reported, potentially assigning multiple themes to each reason.
3.4.3. Control Variables
We assessed participants’ general attitudes toward AI as a control variable in our analysis. This measurement was taken before informing participants about the task.
  • General attitudes towards AI were measured by asking participants to respond to statements that measured their comfort with, perceptions of, and willingness to engage with AI technologies (derived from (Durndell and Haag, 2002) e.g., ”AI is bringing us into a bright new era” and ”AI makes me uncomfortable because I do not understand it”). Responses were recorded using a 7-point Likert scale, where 1 corresponds to ”Strongly disagree” and 7 to ”Strongly agree.”
The Pearson correlation coefficients between the control variable and all dependent variables did not reach statistical significance (all p-values ¿ 0.05)  [^6] Specific data are presented in Appendix 7.2..
4. Results
4.1. Greater Psychological Distance Increases Relative User Perceptions of Conversational Search
Our results generally show that as psychological distance increases from near to far, relative ratings of conversational search (i.e. ConvSearch rating - WebSearch rating) across multiple metrics increased. At near psychological distances, users were generally indifferent between the two systems; at far distances, users instead mostly preferred ConvSearch.
4.1.1. Partial Increase in Relative Perceived Ease of Use (H1 Partially Supported)
Figure 3. Comparison of perceived ease of use between WebSearch and ConvSearch In Far Psychological Distances. Panels (a)-(d) show results at a near psychological distance, and panels (e)-(h) at a far psychological distance, highlighting significant differences in the latter. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
[…]
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived ease of use for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived ease of use at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.490), marked as non-significant (ns); (b) Temporal Distance (p = 0.270), marked as non-significant (ns); (c) Social Distance (p = 0.153), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.560), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived ease of use at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. Panel (e) Spatial Distance shows a p-value of (p = 0.095), marked as non-significant (ns). The p-values indicate significant differences in panels (f) Temporal Distance (p = 0.025), marked with one asterisk (*); (g) Social Distance (p = 0.003), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.004), marked with two asterisks (**).
When psychological distance was near, participants reported similar scores of perceived ease of use for WebSearch (Spatial Distance: $M=6.19$, $SD=0.96$; Temporal Distance: $M=5.91$, $SD=0.79$; Social Distance: $M=5.63$, $SD=1.32$; Hypothetical Distance: $M=5.81$, $SD=1.48$) and ConvSearch (Spatial Distance: $M=5.91$, $SD=1.35$; Temporal Distance: $M=5.55$, $SD=1.22$; Social Distance: $M=6.01$, $SD=0.97$; Hypothetical Distance: $M=6.03$, $SD=1.09$). As depicted in Fig. 3 (a)-(d), paired samples T-Tests showed non-significant differences with low effect sizes (Spatial Distance: $t(15)=0.71$, $p=0.49$, Cohen’s $d=0.18$; Temporal Distance: $t(15)=1.15$, $p=0.27$, Cohen’s $d=0.29$; Social Distance: $t(15)=-1.51$, $p=0.15$, Cohen’s $d=-0.38$; Hypothetical Distance: $t(15)=-0.60$, $p=0.56$, Cohen’s $d=-0.15$).
In contrast, there were significant differences in perceived ease of use at a far psychological distance in three out of the four dimensions (Fig. 3 (f)-(h)). For temporal distance, ConvSearch demonstrated a higher ease of use ($M=6.18$, $SD=0.73$) compared to WebSearch ($M=5.65$, $SD=1.09$), with significant results, $t(15)=-2.14$, $p=\textbf{0.025}$, Cohen’s $d=-0.53$. Similarly, for social distance, ConvSearch ($M=6.22$, $SD=0.74$) outperformed WebSearch ($M=5.26$, $SD=1.10$), $t(15)=-3.14$, $p=\textbf{0.003}$, Cohen’s $d=-0.78$, and for hypothetical distance, ConvSearch ($M=6.30$, $SD=0.76$) was rated higher than WebSearch ($M=5.40$, $SD=1.23$), $t(15)=-3.06$, $p=\textbf{0.004}$, Cohen’s $d=-0.77$. However, with spatial distance (Fig. 3 (e)), although ConvSearch scored higher ($M=6.28$, $SD=0.48$) compared to WebSearch ($M=6.02$, $SD=0.73$), the difference was not statistically significant, $t(15)=-1.37$, $p=0.095$, Cohen’s $d=-0.34$. These findings mostly support H1, indicating that as psychological distance increases, ConvSearch tends to be perceived as easier to use compared to WebSearch, particularly in the temporal, social, and hypothetical dimensions.
4.1.2. Increase in Relative Perceived Information Credibility (H2 Supported)
Figure 4. Impact of psychological distance on perceived information credibility for WebSearch and ConvSearch. Panels (a)-(d) illustrate results at a near psychological distance, with significant differences only in spatial distance. Panels (e)-(h) demonstrate increased credibility for ConvSearch at a far psychological distance across all dimensions except temporal distance. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
[…]
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived information credibility for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived information credibility at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.011), marked with one asterisk (*); (b) Temporal Distance (p = 0.131), marked as non-significant (ns); (c) Social Distance (p = 0.612), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.550), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived information credibility at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.001), marked with three asterisks (***); panels (f) Temporal Distance (p = 0.012), marked with one asterisk (*); (g) Social Distance (p ¡ 0.001), marked with three asterisks (***); (h) Hypothetical Distance (p ¡ 0.001), marked with three asterisks (***).
At a near psychological distance, as depicted in Fig. 4 (b)-(d), participants reported similar perceptions of information credibility in three out of four dimensions between WebSearch (Temporal: $M=76.88$, $SD=15.86$; Social: $M=80.69$, $SD=18.37$; Hypothetical: $M=75.40$, $SD=25.22$) and ConvSearch (Temporal: $M=70.81$, $SD=21.86$; Social: $M=78.88$, $SD=22.58$; Hypothetical: $M=74.40$, $SD=27.80$). These differences were not statistically significant (Temporal: $t(15)=1.60$, $p=0.13$, Cohen’s $d=0.40$; Social: $t(15)=0.52$, $p=0.61$, Cohen’s $d=0.13$; Hypothetical: $t(15)=0.61$, $p=0.55$, Cohen’s $d=0.15$). However, for spatial distance (Fig. 4 (a)), ConvSearch ($M=75.46$, $SD=15.63$) demonstrated a significant credibility advantage over WebSearch ($M=59.28$, $SD=16.38$), $t(15)=-2.90$, $p=0.011$, Cohen’s $d=-0.73$.
In contrast, when psychological distance was far, the perceived credibility of information from ConvSearch was significantly higher compared to WebSearch in all four dimensions (Fig. 4 (e)-(h)). For spatial distance, ConvSearch was rated as more credible ($M=76.38$, $SD=13.28$) than WebSearch ($M=56.88$, $SD=20.03$), yielding a significant difference, $t(15)=-4.92$, $p<\textbf{0.001}$, Cohen’s $d=-1.23$. This was also true for temporal distance (($M=77.94$, $SD=13.36$) vs. ($M=71.21$, $SD=18.73$), $t(15)=-2.50$, $p=\textbf{0.012}$, Cohen’s $d=-0.62$), social distance ($M=85.19$, $SD=13.11$) vs. ($M=70.31$, $SD=20.38$), $t(15)=-3.98$, $p<\textbf{0.001}$, Cohen’s $d=-1.00$), and hypothetical distance ($M=87.54$, $SD=10.11$) vs. ($M=74.23$, $SD=11.37$), $t(15)=-4.12$, $p<\textbf{0.001}$, Cohen’s $d=-1.03$. These findings support H2. They suggest that as psychological distance increases, so does the perceived reliability of information from ConvSearch compared to WebSearch.
4.1.3. Increase in Relative Perceived Usefulness (H3 Supported)
Figure 5. Comparative analysis of perceived usefulness between WebSearch and ConvSearch at varying psychological distances. Panels (a)-(d) show minimal differences at a near psychological distance, except for social distance. Panels (e)-(h) depict significant increases in perceived usefulness for ConvSearch at a far psychological distance across all dimensions. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
[…]
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived usefulness for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived usefulness at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.337), marked as non-significant (ns); (b) Temporal Distance (p = 0.180), marked as non-significant (ns); (c) Social Distance (p = 0.031), marked with one asterisk (*); (d) Hypothetical Distance (p = 0.418), marked as non-significant (ns). In the second row, panels (e) through (h) illustrate the perceived usefulness at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.005), marked with two asterisks (**); panels (f) Temporal Distance (p ¡ 0.001), marked with three asterisks (***); (g) Social Distance (p = 0.001), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.003), marked with two asterisks (**).
At a near psychological distance, perceived usefulness was similar in three out of the four dimensions (Fig. 5 (a),(b) and (d)) between WebSearch (Spatial: $M=6.14$, $SD=1.08$; Temporal: $M=5.87$, $SD=1.09$; Hypothetical: $M=5.44$, $SD=1.79$) and ConvSearch (Spatial: $M=5.68$, $SD=1.50$; Temporal: $M=5.49$, $SD=1.59$; Hypothetical: $M=5.90$, $SD=1.48$). These differences were not significant (Spatial: $t(15)=1.07$, $p=0.300$, Cohen’s $d=0.27$; Temporal: $t(15)=1.35$, $p=0.198$, Cohen’s $d=0.34$; Hypothetical: $t(15)=-0.86$, $p=0.405$, Cohen’s $d=-0.21$). However, with social distance, a significantly higher score ($t(15)=-2.38$, $p=0.031$, Cohen’s $d=-0.60$) for ConvSearch ($M=6.04$, $SD=1.04$) was observed compared to WebSearch ($M=5.33$, $SD=1.50$), as shown in Fig. 5 (c).
In contrast, at a far psychological distance, ConvSearch was perceived as significantly more useful in all four dimensions (Fig. 5 (e)-(h)). With spatial distance, ConvSearch ($M=6.14$, $SD=0.57$) was perceived as more useful compared to WebSearch ($M=5.04$, $SD=1.48$), $t(15)=-2.83$, $p=\textbf{0.006}$, Cohen’s $d=-0.71$. Similar phenomena were observed in the temporal ($M=6.63$, $SD=0.64$ vs. $M=5.26$, $SD=1.41$, $t(15)=-3.79$, $p<\textbf{0.001}$, Cohen’s $d=-0.95$), social ($M=6.54$, $SD=0.59$ vs. $M=4.85$, $SD=1.57$, $t(15)=-3.66$, $p=\textbf{0.001}$, Cohen’s $d=-0.92$), and hypothetical ($M=6.46$, $SD=0.90$ vs. $M=4.81$, $SD=1.78$, $t(15)=-3.29$, $p=\textbf{0.002}$, Cohen’s $d=-0.82$) dimensions. Notably, the social distance dimension not only maintained its significance, but also showed an increased effect size, moving from a marginal to a highly significant difference. These findings support H3, demonstrating that as psychological distance increases, so does the perceived usefulness of ConvSearch compared to WebSearch.
4.1.4. Increase in Relative Perceived Enjoyment (H4 Supported)
Figure 6. Comparative analysis of perceived enjoyment between WebSearch and ConvSearch across four psychological distances. Panels (a)-(c) indicate no significant differences at a near psychological distance, while panel (d) shows a marginal preference for ConvSearch. Panels (e)-(h) demonstrate significant increases in perceived enjoyment for ConvSearch at a far psychological distance across all dimensions. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
[…]
This figure presents eight violin plots arranged in two rows and four columns, comparing the perceived enjoyment for two search technologies, WebSearch and ConvSearch, across four dimensions of psychological distance: spatial, temporal, social, and hypothetical. Each plot features a pair of violin plots side by side, with the left representing WebSearch and the right representing ConvSearch. The violin plots illustrate the distribution of responses, with a central box plot highlighting the median (red line), the interquartile range (colored box), and outliers (diamond shapes). Statistical significance is indicated with p-values and asterisks of the paired samples t-test. In the first row, panels (a) through (d) depict the perceived enjoyment at a near psychological distance. The violin plots are colored in shades of beige and brown, representing WebSearch and ConvSearch respectively. The p-values are as follows: (a) Spatial Distance (p = 0.450), marked as non-significant (ns); (b) Temporal Distance (p = 0.669), marked as non-significant (ns); (c) Social Distance (p = 0.212), marked as non-significant (ns); (d) Hypothetical Distance (p = 0.018), marked with one asterisk (*). In the second row, panels (e) through (h) illustrate the perceived enjoyment at a far psychological distance. The violin plots are colored in shades of green and blue, again representing WebSearch and ConvSearch respectively. The p-values indicate significant differences in panel (e) Spatial Distance (p = 0.014), marked with one asterisk (*); panels (f) Temporal Distance (p ¡ 0.001), marked with three asterisks (***); (g) Social Distance (p = 0.004), marked with two asterisks (**); (h) Hypothetical Distance (p = 0.001), marked with two asterisks (**).
At a near psychological distance, perceived enjoyment was similar in three out of the four dimensions (Fig. 6 (a)-(c)) between WebSearch (Spatial: $M=5.40$, $SD=1.03$, Temporal: $M=5.13$, $SD=1.28$, Social: $M=5.33$, $SD=1.34$) and ConvSearch (Spatial: $M=5.00$, $SD=1.81$, Temporal: $M=4.98$, $SD=1.78$), Social: $M=5.75$, $SD=1.21$). These differences were not significant (Spatial: $t(15)=0.78$, $p=0.450$, Cohen’s $d=0.19$; Temporal: $t(15)=0.44$, $p=0.669$, Cohen’s $d=0.11$; Social: $t(15)=-1.30$, $p=0.212$, Cohen’s $d=-0.33$). However, in the hypothetical distance dimension there was a significant difference (ConvSearch: ($M=5.67$, $SD=1.22$; WebSearch: ($M=4.71$, $SD=1.50$, ($t(15)=-2.65$, $p=0.018$, Cohen’s $d=-0.66$). suggesting a higher perceived enjoyment of ConvSearch even at a near psychological distance (Fig. 6 (d)).
In contrast, when psychological distance was far, there were significant differences in all four dimensions (Fig. 6 (e)-(h)). ConvSearch produced greater enjoyment than WebSearch (Spatial: $M=5.46$, $SD=1.05$ vs. $M=4.67$,$SD=1.05$,$t(15)=-2.42$, $p=\textbf{0.014}$, Cohen’s $d=-0.60$; Temporal: $M=5.92$, $SD=1.23$ vs. $M=4.73$, $SD=1.41$, $t(15)=-4.73$, $p<\textbf{0.001}$, Cohen’s $d=-1.18$; Social: $M=5.65$, $SD=1.21$ vs. $M=5.65$, $SD=1.36$, $t(15)=-3.03$, $p=\textbf{0.004}$, Cohen’s $d=-0.76$; Hypothetical: $M=6.00$, $SD=1.23$ vs. $M=4.63$, $SD=1.49$, $t(15)=-3.67$, $p=\textbf{0.001}$, Cohen’s $d=-0.92$). Notably, the difference in perceived enjoyment in the hypothetical distance was greater in the far condition compared to the near condition. These results underscore that as psychological distance increases, the perceived enjoyment of using ConvSearch significantly exceeds that of WebSearch across various dimensions, supporting H4.
4.1.5. Increased Preferences for Conversational Search (H5 Supported)
Figure 7. Differential preference for ConvSearch over WebSearch across various psychological distances, shown in two parts: (a) detailed comparison across four dimensions, and (b) aggregated preference at near and far distances. Significance is marked as p ¿= 0.05 (ns), p ¡ 0.05 (*), p ¡ 0.01 (**), or p ¡ 0.001 (***).
[…]
This figure is divided into two parts labeled as (a) and (b), each depicting the preference for ConvSearch over WebSearch across different psychological distances. Part (a) features a horizontal bar chart with eight bars grouped into four pairs, each representing a different psychological dimension: Spatial, Temporal, Social, and Hypothetical. Each pair consists of a beige bar for ’Near’ psychological distance and a purple bar for ’Far’ psychological distance. The y-axis is scaled from 0 to 1 (0% to 100%) and is labeled ”Preference for ConvSearch.” The exact percentages of preference for ConvSearch are displayed above each bar. For Spatial Distance, the beige bar shows 75.0% and the purple bar shows 87.5%. For Temporal Distance, the beige bar is at 56.2% and the purple bar at 87.5%. In Social Distance, the beige bar indicates 75.0% and the purple bar 81.2%. Finally, for Hypothetical Distance, the beige bar is at 75.0% and the purple bar significantly higher at 93.8%. Part (b) of the figure presents a simpler bar chart with just two bars, again in beige and purple, representing the aggregated preference for ConvSearch at Near and Far psychological distances, respectively. The beige bar shows a preference of 71.9%, while the purple bar shows a higher preference of 89.1%. The significance of the difference is indicated by an asterisk and the note (p = 0.026) above the bars, referring to the results of a chi-square test.
The impact of psychological distance on the acceptance of ConvSearch supports H5. When the psychological distance is close (as shown in Fig. 7 (a)), the preference for ConvSearch is already evident, with 75.0% of participants favoring it in the spatial dimension. However, as the distance increases, this preference increases to 87.5%. Similar patterns are observed in other dimensions as psychological distance increases: for temporal distance, the preference increases from 56.2% social distance, from 75.0% to 81.2%; and in hypothetical distance,93.8%. Upon aggregating the data across all dimensions and analyzing the preference for WebSearch and ConvSearch at near and far psychological distances (Fig. 7 (b)), a chi-square test revealed significant differences ( $\chi^{2}(1,N=128)=4.97$, $p=\textbf{0.026}$). Thus, while participants exhibit a preference for ConvSearch even at a near psychological distance, this preference intensifies significantly at a far psychological distance.
4.2. Thematic Analysis of User Preferences
We also analysed the qualitative responses regarding participants’ preferences for one system over the other. Some of the themes that emerged were unique to a particular system (i.e. only participants who preferred that system mentioned that theme), while others were shared between both systems.
4.2.1. Unique Themes for Preferring ConvSearch
These themes highlight the distinct advantages of ConvSearch over WebSearch.
Low Workload. Many participants (n = 58) reported that using ConvSearch required less effort compared to WebSearch. They felt that ConvSearch simplified the information gathering process by eliminating the need to visit multiple websites and compile information independently. As one participant (P6) explained: “It was easier to find out what I needed to know and less work on my part to pull it all together.” Thus, reduced workload was one of the reasons for their preference for ConvSearch.
Speed and Efficiency. Participants (n = 43) perceived ConvSearch as faster and more efficient in making travel plans. The efficiency they focused on was whether ConvSearch could quickly provide information while meeting their needs to complete travel plans. As P108 stated, “I would choose the chatbot because it saves me time and is more efficient at making a full-bodied itinerary.
Specific Information Consolidation. Another major reason participants (n = 43) preferred ConvSearch was its ability to quickly consolidate information from various sources for specific queries. P96 said, “The results the chatbot provides is neater compared to the websearch’s results which are just links while the chatbot makes sure to summarize any relevant info and provides it to me.
Interactive and Conversational. ConvSearch’s natural dialogue offers participants (n = 18) a superior interactive experience. Some participants described the interaction as akin to conversing with a human. For instance: “I would love to use the chatbot since it feels like having a conversation with a human and getting personalized advice from them. This makes the information easier to understand and apply to my travel plans.” (P77)
4.2.2. Unique Themes for Preferring WebSearch
WebSearch has fewer unique themes than ConvSearch, but nevertheless has advantages in certain areas.
Visual Confirmation and Source Credibility. Some participants (n = 12) mentioned that WebSearch provides information not only in text but also in images. In contrast to ConvSearch, WebSearch enables participants to identify the sources of information. This capability allows participants to evaluate the reliability of the information. The ability to visually confirm and select information sources plays a significant role in users’ decision-making processes. One participant called it, “I like to see what I am getting into and have a visual to confirm that I am making a good choice. It makes me feel more at ease with my decision.” (P2)
Specifically, for time-sensitive information, WebSearch is uniquely able to help users verify timeliness. As one participant commented: “I’d rather use the web search so I… can see how recent the information is.” (P121)
Habit and Experience. A minority of participants (n = 2) prefer WebSearch due to habitual use and greater experience with it. One participant called it, “I think that I can access more information and feel more confident about where that information is coming from on a web search. I’ve been doing web searches forever and feel confident in my ability to find the best and most credible information through them and I can bookmark my findings for later.” (P10)
4.2.3. Themes Found in Both Systems
Some themes emerged in participants’ reasons for preferring both ConvSearch and WebSearch. These themes generally reflect subjective perceptions rather than functional differences between ConvSearch and WebSearch, indicating personalized preferences.
Reliability and Accuracy. Some participants (n = 7) cited the inaccuracy and unreliability of ConvSearch as reasons for preferring WebSearch. As one participant explained: “I would use the web search, although it may be more cumbersome. The reason is that I am able to see the direct sources for the information I receive, unlike the chatbot. The chatbot was easier, just not reliable enough for me.” (P80)
Others (n = 12) expressed contrary views, believing that web search results might contain more misinformation, making conversational search more reliable and accurate. One noted: “I would choose the chatbot rather than a web search because I would not have to look through multiple web pages for information and potentially run into misinformation and instead I can get trustworthy answers from the chatbot itself.” (P117)
Discovery and Exploration. WebSearch’s result list offered participants (n = 4) greater freedom to explore information and gain new ideas and interests during the exploration process. One participant said: “I would rather use the web search because I like to look things up and decide which sites I want to use for information on my own. Also, when you do a web search you may find things that spark new ideas and interests. I think it might be useful to use a chatbot to get quick bits of information, and the use the web to search for the things you are most interested in.” (P18)
For some participants (n = 9), the more comprehensive information provided by ConvSearch served as a foundation to gain new ideas and further explore the information. One participant said: “I think it’s a good foundational point to work off of, especially when creating an itinerary at a location I know nothing about. Chatbot gave me ideas of what to explore further and potentially search to see reviews before finalizing.” (P116)
Combining Two Tools  [^7] Detailed data can be found in the Appendix 7.3.. Some participants (n = 9), despite being explicitly instructed to provide reasons for their preference for either web search or conversational search, still suggested combining both tools. As one noted: “Chatbot makes it easier to get a lot of info as a great starting point. I would then use a search engine to verify the info from the chatbot.” (P32)
Comprehensive Information. For planning trips, one participant (n = 1) preferred to refer to real experiences shared by other travelers. WebSearch can provide these authentic experiences, which ConvSearch cannot. Therefore, WebSearch offers more comprehensive information. “I would prefer a search box because it gives more details and articles from previous experiences.” (P43)
Participants (n = 7) who believed ConvSearch provided more information did so because it delivered more content than WebSearch for the same amount of effort or time. For instance: “I think the chatbot could give me more information depending on the prompt I gave it.” (P4)
5. Discussion
Our findings highlight the pivotal role of psychological distance of information seeking tasks in shaping user perceptions and preferences toward different IR technologies. As psychological distance increased, participants showed a preference for LLM-powered conversational search over web search, perceiving it as easier to use, more credible, useful, and enjoyable. Consequently, user acceptance of LLM-powered conversational search also rose with increasing psychological distance. These critical insights highlight how the ubiquitous factor of psychological distance informs design considerations for the nascent technology that is generative AI-based conversational search.
5.1. Deep Impact of Psychological Distance in Information Seeking Tasks (RQ)
Our findings corroborate CLT’s suggested existence of a relationship between psychological distance and user evaluations, and hence choice of IR technologies (Liberman and Trope, 2014; Trope and Liberman, 2010; Trope et al., 2007). People’s choices and preferences are based on their understanding of events, which is in turn affected by psychological distance (Trope and Liberman, 2010). Our research provides a comprehensive view from the nature of tasks to user preferences, not only enhancing our understanding of the effects of task psychological distance but also offering practical guidance for future IR system design. By addressing user information needs at different psychological distances, IR systems can further improve user interaction experiences and satisfaction.
5.1.1. Greater Psychological Distance Makes Conversational Search Easier to Use (H1)
Our results suggest that tasks with greater psychological distance make participants find conversational search easier than web search. This is due to conversational search providing integrated and summarized information tailored to specific questions (Kaushik and Jones, 2023). It aligns better with the high-level abstract thinking used by users in these situations (Hadar et al., 2022; Yoo and Lee, 2023). Conversely, web search provides a list of results, requiring users to filter and summarize information themselves.
Our thematic analysis supports this perspective. Participants cited themes Low Workload and Specific Information Consolidation for their preference for conversational search.Low Workload is linked to its perceived ease of use, as less workload allows users to meet their needs more easily. Specific Information Consolidation is seen as a reason for Low Workload, illustrating how conversational search’s ability to integrate information underpins its ease of use.
5.1.2. Greater Psychological Distance Makes Conversational Search More Credible (H2)
Our results show that tasks with greater psychological distance lead participants to perceive higher information credibility in LLM-powered conversational search than web search. This perception is further elucidated by our thematic analysis. In the themes reported by participants, Reliability and Accuracy are closely related to their perception of information credibility. Participants who prefer conversational search believe it provides more accurate answers by helping them filter results. In contrast, web search results often provide too many options, which sometimes do not align with the information sought by users.
Further analysis of the three dimensions of credibility  [^8] Specific data are presented in Appendix 7.4. reveals that although user perceptions of the accuracy and believability of information from LLM-powered conversational search significantly improve with greater psychological distance, the perceived enhancement in information authenticity is not as pronounced. This may be because assessing information authenticity depends more on directly verifying information sources, an area where LLM-powered conversational search is weaker compared to web search (Bender et al., 2021). Additionally, conversational search is limited by its LLM core, which is prone to hallucinations (Maynez et al., 2020) and relies on outdated training data (Achiam et al., 2023), further reducing user perception of the authenticity of the information provided. Therefore, LLM-powered conversational search still needs to strengthen its capabilities regarding verifiable information sources, hallucination issues, and providing up-to-date information to address concerns about information authenticity.
5.1.3. Greater Psychological Distance Makes Conversational Search More Useful and Enjoyable (H3-H4)
As psychological distance is far, user perceptions of the usefulness and enjoyment of LLM-powered conversational search improve significantly. This underscores that increasing the psychological distance of a task can enhance user perceptions of the usefulness and enjoyment of conversational search, even without changes in the technology itself. This finding aligns with the TAM and its extensions (Davis, 1989; Davis et al., 1992, 1992), which shows the positive interlink between perceived ease of use (Adams et al., 1992), credibility (Ayeh, 2015), enjoyment, and usefulness. In our thematic analysis, Speed and Efficiency are directly related to user perception of usefulness. Participants find conversational search useful because it allows them to complete tasks faster and more efficiently than web search. The perception of enjoyment of conversational search results from enhanced interactivity through natural dialogue (theme Interactive and Conversational). Some participants feel that interacting with conversational systems is like conversing with another human, aligning with research findings that humanlike qualities are a significant factor in making chatbots enjoyable (Følstad and Brandtzaeg, 2020).
This finding is particularly noteworthy as it suggests a link between two seemingly unconnected topics - psychological distance of a search task and enjoyment using a search system. Psychological distance and enjoyment have been studied in areas such as teaching (Büssing et al., 2020) and sports watching (Lim et al., 2012), where user enjoyment is naturally paramount to user outcomes. However, the fact that this relation persists even in a task like IR - not something intuitively perceived as inducing a high level of enjoyment - suggests that IR system designers should consider psychological distance as an important factor in determining user satisfaction and outcomes.
5.2. Overall User Acceptance of Conversational Search (H5)
Our findings align with the TAM and its extensions (Davis, 1989; Davis et al., 1992), suggesting that improvements in the previous four user metrics culminate in increased user acceptance of LLM-powered IR systems as psychological distances increase (H5). Additionally, we found that controlling for psychological distance, users generally prefer LLM-powered conversational search over traditional web search. For common information seeking tasks like travel planning, this preference may stem from the inherent advantages of LLM-powered conversational search in areas such as Low Workload, Speed and Efficiency, Specific Information Consolidation, and Interactive and Conversational capabilities. These characteristics enable LLM-powered conversational search systems to more effectively meet users’ information needs in daily tasks.
5.3. Design Implications
Our research confirms the significant impact of the psychological distance of information seeking tasks on user perception and preferences. Conversational search, with its ability to provide information from an abstract perspective, better meets user needs when the psychological distance is greater. However, current conversational search systems exhibit a fixed level of abstraction in information provision, regardless of human cognitive state (Kirshner, 2024). This limits the effectiveness of LLMs in providing information, as it does not fully meet the diverse judgment needs inherent in human cognition (Santurkar et al., 2023). Past HCI research has also touched on similar views when discussing the design of conversational search. They identified user demands for different levels of information, with some preferring direct access to answers, while others favored receiving broad background information (Vtyurina et al., 2017). This aligns with the two information levels that we have elucidated. However, these studies did not identify the key factors behind these differing information needs.
We propose using psychological distance as a primary guiding factor to enhance the diversity of information levels provided by conversational search systems. By considering psychological distance, we can not only increase the diversity of information, but also better meet the information needs at the user’s current cognitive state.
5.3.1. Impact on Conversational Search Design
In further enhancing LLM-powered conversational search systems, it is crucial to fully leverage the capabilities of LLMs by assessing the psychological distance of information seeking tasks through dialogue content and treating it as a key attribute. For example, the system could evaluate the psychological distance of tasks by analyzing user descriptions across the spatial, temporal, social, and hypothetical dimensions. Additionally, this assessment could use the capability of conversational search to pose supplementary questions (Radlinski and Craswell, 2017), refining the evaluation of task psychological distance through additional inquiries focused on these dimensions.
Conversational search systems could adjust their IR strategies based on the psychological distance perceived by users. For tasks with a closer psychological distance, users typically seek specific and actionable information. In such cases, search results should focus on providing detailed steps and practical guidance. For example, when planning a trip in the near future, the system could provide specific travel itineraries, recommended accommodations, and dining options. This approach not only satisfies the user’s low-level, specific information needs but also aids in making quick decisions and executing plans.
Conversely, for tasks with greater psychological distance, users tend to seek higher-level, more abstract information. Here, conversational search could provide an overview of the topic, helping users quickly grasp the core content and related broad information. For instance, if a user is considering potential future travel destinations, the system could provide summaries of various cities, their cultural highlights, and culinary recommendations, thus providing a comprehensive understanding and sparking further exploration interest.
5.3.2. Combine Conversational Search and Web Search
Web search, a widely used IR technology, offers significant advantages in visualization and source reliability that should not be overlooked. Currently, clear guidance on integrating conversational search with web search is lacking, often leaving users to choose their preferred tool without knowing which would better satisfy their information needs. We suggest offering tools that more closely align with current information needs based on the psychological distance of a given task. Therefore, we propose several recommendations for systems that integrate conversational and web search.
Initially, it would be beneficial to present a conversational search interface to users. This design leverages the natural conversational abilities of LLMs to reduce the effort users must expend in formulating queries (Radlinski and Craswell, 2017). The system can utilize the LLM’s understanding capabilities to assess the psychological distance of the user’s current task through dialogue and to ask supplementary questions (Radlinski and Craswell, 2017), thereby more accurately gathering information about the task’s psychological dimensions. This step can enable the system to tailor the presentation of information more effectively.
For tasks with closer psychological distances, we recommend returning a list of web search results. This is because when the task’s psychological distance is close, users typically need specific and detailed information, and web search can provide a rich and detailed list of sources. For example, when searching for local restaurants or immediate weather forecasts, web search can provide timely and specific information that directly meets users’ needs.
However, for tasks with a greater psychological distance, we suggest providing synthesized information generated by conversational search. In these cases, users tend to seek overview or conceptual information, and conversational search can meet these needs by integrating and summarizing information. For example, when exploring potential travel destinations or learning about new technologies, conversational search can provide a comprehensive information framework to help users build a preliminary understanding of the topic.
Although past HCI research has not focused on the psychological distance of tasks, it has explored factors influencing user choices between conversational and web search in specific task scenarios. A study on programming tasks noted that programmers opt for conversational search when objectives are clear, seeking results aligned with their goals (Yen et al., 2024). This phenomenon, stemming from the confirmatory nature of conversational search, known as the echo chamber effect, has been extensively studied in HCI (Sharma et al., 2024; Tanprasert et al., 2024). This characteristic of conversational search limits users’ exposure to diverse viewpoints and is difficult to fundamentally resolve due to its LLM core. To optimize this, some studies have suggested designs that allow conversational systems to actively present different viewpoints (Feltwell et al., 2020) and designs that proactively remind users of current information biases in conversations (Sharma et al., 2024). However, our proposed solution is to integrate conversational and web search, leveraging the diversity of viewpoints offered by web search to balance the confirmatory bias of conversational search. Additionally, HCI research discussing user trust in ChatGPT versus web search noted that generative AI’s lack of information sources reduces user trust (Jung et al., 2024). Yet, its assistance in ideation encourages its use. Our participants expressed similar views [^9] Specific data are presented in Appendix 7.3., preferring to obtain preliminary information through conversational search, then verify this information and acquire more details through web search.
Therefore, we propose a deeper integration of the two tools. Even when providing synthesized information from conversational search, users should be allowed to click on keywords or content mentioned in the synthesis to directly transition to web search. Through this approach, our goal is to use the web search to help users verify the accuracy of related information, while retaining the benefits of conversational search.
5.4. Limitations
This study has some limitations. First, our study only considered information seeking tasks in travel scenarios. These tasks, being general in nature, may have influenced user perceptions of LLM-powered conversational search, which is better suited for general rather than specialized tasks. This limitation might obscure the specific differences in user perceptions between web search and LLM-powered conversational search. For instance, we observed no significant differences in most metrics when the psychological distance was near. This pattern may not apply to other scenarios and also prevented us from further exploring the specific impacts of tasks with near psychological distance on both types of IR tools. However, with web search as a control, we found no significant differences in most ratings between the two IR tools when the psychological distance was near, but significant differences emerged when the psychological distance was far. This outcome still addresses our main research question and reveals a general principle: Users prefer LLM-powered conversational search when the psychological distance is greater.
Furthermore, the custom search interface we implemented does not fully represent a real, commercial search page; it lacks features like image previews and related searches, especially compared to the actual Google Search website. However, since the search results are sourced from the genuine Google Search API, we consider them to still fairly represent real search interactions.
6. Conclusion
In this study, we employed a mixed-methods experiment to explore how psychological distance in information seeking tasks affects user perceptions and preferences for IR technologies. We found that user preference for conversational search increases with the psychological distance of the task, and that this preference is primarily driven by users’ perceptions of ease of use, credibility, usefulness, and enjoyment. Notably, these perceptions are significantly enhanced in conversational search when the task’s psychological distance is greater. By integrating CLT with IR technologies, our study underscores the psychological distance in tasks as a key factor that influences user perceptions and shapes user preferences. conversational systems to adjust their output based on task nature, and for conversation-web search hybrid systems that take into account the nature of different tasks.
7. Appendix
7.1. Task Scenarios for Different Dimensions of Psychological Distance
Figure 8. Overview of task scenarios used in the study, depicting variations in travel plans across different psychological distances and dimensions. Each scenario is tailored to manipulate spatial, temporal, social, and hypothetical distances to explore their impact on user perceptions and technology preference.
[…]
This figure presents a table with four rows and two columns, each row representing a different psychological dimension: spatial, temporal, social, and hypothetical. The first column lists the dimension and the psychological distance (near or far), and the second column provides detailed descriptions of the travel scenarios used in the study. For spatial distance, near scenarios involve travel to Mexico or Cuba next month, while far scenarios involve travel to Sudan or Egypt next month. For temporal distance, near scenarios describe travel next week, and far scenarios next year, both to Mexico or Cuba. Social distance scenarios vary by the relationship to the vacationer, either a close friend or an internet user. Hypothetical distance scenarios differ in the certainty of the trip, with near scenarios having confirmed plane tickets and far scenarios being unconfirmed. Each scenario is designed to assess how psychological distance influences user interaction with WebSearch and ConvSearch technologies.
Figure 8 provides an extensive description of the task scenarios used in the study, which are designed to manipulate psychological distance across four dimensions: spatial, temporal, social, and hypothetical. Each scenario is crafted to maintain consistency in activities and destinations while varying only the dimension of psychological distance under investigation. This design ensures that the effects observed can be attributed specifically to changes in psychological distance rather than other factors.
7.2. Correlations Between Control Variables and Dependent Variables
We conducted Pearson correlation analyses to explore the relationships between the general attitude towards AI and various dependent variables related to user interactions with web search (WebSearch) and conversational search (ConvSearch) systems. The results of these analyses are summarized in the Table 3:
Table 2. Pearson Correlations between General Attitude towards AI (Control Variable) and Dependent Variables
Dependent Variables $r$ $p$
Perceived Ease of Use of WebSearch -0.033 0.708
Perceived Information Credibility of WebSearch 0.129 0.148
Perceived Usefulness of WebSearch 0.015 0.863
Perceived Enjoyment of WebSearch 0.030 0.735
Perceived Ease of Use of ConvSearch 0.050 0.578
Perceived Information Credibility of ConvSearch -0.014 0.876
Perceived Usefulness of ConvSearch 0.068 0.445
Perceived Enjoyment of ConvSearch 0.080 0.371
User Preference -0.017 0.850
Table 3. Note: All $p$-values are two-tailed. Pearson correlation coefficients ($r$) measure the strength and direction of the linear relationship between the general attitude towards AI (as a control variable) and main dependent variables.
7.3. Participant Responses About User Preferences Reasons of Combining Two Tools
In our study, a notable subset of participants expressed a preference for combining the functionalities of ConvSearch and WebSearch. Despite being instructed to choose between the two, these participants highlighted the complementary benefits of using both tools in tandem. Below are detailed reasons provided by participants for preferring a combined use of ConvSearch and WebSearch:
  • “I would give the chatbot a chance but would confirm its plan with traditional searches.” (P9)
  • “For this type of ’big picture’ task, chatbots are great for coming up with ideas. I would start here and then move to search to find out specifics and validate the chatbot’s information.” (P24)
  • “Chatbot makes it easier to get a lot of info as a great starting point. I would then use a search engine to verify the info from the chatbot.” (P32)
  • “I would probably use the chatbot to gain preliminary information, but then double check information using a search engine.” (P38)
  • “I would use the chatbot to plan it out, but would use a web search to make sure the information is correct. So I would realistically use a combination of the two, just to check its suggestions.” (P40)
  • “Although both search engine and chatbot helped me to achieve my goals, I believe web search can offer more detailed descriptions of my searches. Chatbot is accurate but does not provide as much information in one search as search engine does.” (P66)
  • “I like that the chatbot can come up with some good ideas and put together an entire plan quickly and then I can look around online for more details if needed to confirm things and book them.” (P69)
  • “I probably would start with a chatbot to get some general ideas. Once I have more information I would probably use a web search to get more detail (independent) information.” (P81)
  • “I think chatbot would be faster but I would have more control over what I want using web search and I would know the source of the information as well.” (P126)
7.4. Results of User Perceptions on Information Accuracy, Authenticity, and Believability
7.4.1. Results of User Perceptions on Information Accuracy
At a near psychological distance, as illustrated in Fig. 9 (a)-(d), there were no significant differences in the perceived information accuracy between WebSearch and ConvSearch across spatial, temporal, social, and hypothetical distances. Specifically, for spatial distance, WebSearch ($M=83.81$, $SD=14.58$) and ConvSearch ($M=77.69$, $SD=14.29$) showed no significant difference ($t(15)=1.57$, $p=0.138$, Cohen’s $d=0.39$). Similar patterns were observed in temporal distance (WebSearch $M=75.75$, $SD=14.44$; ConvSearch $M=70.63$, $SD=19.34$; $t(15)=1.13$, $p=0.278$, Cohen’s $d=0.28$), social distance (WebSearch $M=81.50$, $SD=19.43$; ConvSearch $M=79.38$, $SD=23.09$; $t(15)=0.61$, $p=0.554$, Cohen’s $d=0.15$), and hypothetical distance (WebSearch $M=77.19$, $SD=22.04$; ConvSearch $M=74.06$, $SD=26.65$; $t(15)=1.00$, $p=0.335$, Cohen’s $d=0.25$).
Figure 9. Comparative analysis of perceived information accuracy between WebSearch and ConvSearch at varying psychological distances.
[…]
However, at a far psychological distance, as shown in Fig. 9 (e)-(h), significant differences emerged. In spatial distance, ConvSearch ($M=75.38$, $SD=14.56$) was rated significantly higher in accuracy than WebSearch ($M=69.50$, $SD=18.03$), yielding a significant difference ($t(15)=-2.13$, $p=0.025$, Cohen’s $d=-0.53$). For temporal distance, ConvSearch ($M=74.13$, $SD=15.02$) also demonstrated higher accuracy compared to WebSearch ($M=67.13$, $SD=19.83$), with significant results ($t(15)=-2.15$, $p=0.024$, Cohen’s $d=-0.54$). In the social distance dimension, ConvSearch ($M=84.88$, $SD=13.88$) was perceived as more accurate than WebSearch ($M=66.25$, $SD=21.45$), $t(15)=-4.83$, $p<0.001$, Cohen’s $d=-1.21$. Lastly, for hypothetical distance, ConvSearch ($M=87.25$, $SD=10.64$) significantly outperformed WebSearch ($M=73.81$, $SD=9.88$), $t(15)=-3.87$, $p<0.001$, Cohen’s $d=-0.97$.
7.4.2. Results of User Perceptions on Information Authenticity
At a near psychological distance, as depicted in Fig. 10 (a)-(d), there were no significant differences in the perceived information authenticity between WebSearch and ConvSearch across various dimensions. Specifically, for spatial distance, WebSearch ($M=80.00$, $SD=17.62$) and ConvSearch ($M=71.88$, $SD=20.76$) showed no significant difference ($t(15)=1.52$, $p=0.149$, Cohen’s $d=0.38$). Similar patterns were observed in temporal distance (WebSearch $M=74.38$, $SD=22.71$; ConvSearch $M=66.25$, $SD=27.47$; $t(15)=1.80$, $p=0.092$, Cohen’s $d=0.45$), social distance (WebSearch $M=78.06$, $SD=20.34$; ConvSearch $M=77.19$, $SD=25.54$; $t(15)=0.20$, $p=0.843$, Cohen’s $d=0.05$), and hypothetical distance (WebSearch $M=70.19$, $SD=29.68$; ConvSearch $M=71.75$, $SD=35.19$; $t(15)=-0.70$, $p=0.498$, Cohen’s $d=-0.17$).
Figure 10. Comparative analysis of perceived information authenticity between WebSearch and ConvSearch at varying psychological distances.
[…]
However, at a far psychological distance, as shown in Fig. 10 (e)-(h), significant differences emerged in some dimensions. In spatial distance, there was no significant difference between ConvSearch ($M=73.44$, $SD=14.27$) and WebSearch ($M=69.25$, $SD=21.08$), with $t(15)=-1.51$, $p=0.076$, Cohen’s $d=-0.38$. For temporal distance, no significant difference was found between ConvSearch ($M=74.94$, $SD=17.77$) and WebSearch ($M=69.44$, $SD=20.18$), $t(15)=-1.28$, $p=0.111$, Cohen’s $d=-0.32$. However, in the social distance dimension, ConvSearch ($M=83.38$, $SD=14.18$) was perceived as significantly more authentic than WebSearch ($M=71.31$, $SD=20.35$), $t(15)=-3.07$, $p=0.004$, Cohen’s $d=-0.77$. Lastly, for hypothetical distance, ConvSearch ($M=85.88$, $SD=11.92$) significantly outperformed WebSearch ($M=73.81$, $SD=15.12$), $t(15)=-3.42$, $p=0.002$, Cohen’s $d=-0.86$.
7.4.3. Results of User Perceptions on Information Believability
At a near psychological distance, as illustrated in Fig. 11 (a)-(d), there were no significant differences in the perceived information believability between WebSearch and ConvSearch across various dimensions. Specifically, for spatial distance, WebSearch ($M=85.25$, $SD=14.60$) and ConvSearch ($M=76.81$, $SD=18.52$) showed no significant difference ($t(15)=1.48$, $p=0.159$, Cohen’s $d=0.37$). Similar patterns were observed in temporal distance (WebSearch $M=75.56$, $SD=21.56$; ConvSearch $M=80.50$, $SD=13.94$; $t(15)=1.15$, $p=0.268$, Cohen’s $d=0.287$), social distance (WebSearch $M=82.50$, $SD=17.52$; ConvSearch $M=80.06$, $SD=20.99$; $t(15)=0.77$, $p=0.456$, Cohen’s $d=0.19$), and hypothetical distance (WebSearch $M=77.38$, $SD=26.53$; ConvSearch $M=78.81$, $SD=25.57$; $t(15)=0.72$, $p=0.485$, Cohen’s $d=0.18$).
Figure 11. Comparative analysis of perceived information believability between WebSearch and ConvSearch at varying psychological distances.
[…]
However, at a far psychological distance, as shown in Fig. 11 (e)-(h), significant differences emerged in some dimensions. In spatial distance, ConvSearch ($M=80.31$, $SD=14.39$) was perceived as significantly more believable than WebSearch ($M=73.25$, $SD=21.27$), with $t(15)=-2.02$, $p=0.031$, Cohen’s $d=-0.51$. For temporal distance, ConvSearch ($M=84.75$, $SD=14.13$) significantly outperformed WebSearch ($M=77.06$, $SD=21.79$), $t(15)=-1.98$, $p=0.033$, Cohen’s $d=-0.49$. In the social distance dimension, ConvSearch ($M=87.31$, $SD=12.32$) was perceived as significantly more believable than WebSearch ($M=73.38$, $SD=20.98$), $t(15)=-3.33$, $p=0.002$, Cohen’s $d=-0.83$. Lastly, for hypothetical distance, ConvSearch ($M=89.50$, $SD=9.63$) significantly outperformed WebSearch ($M=75.06$, $SD=12.73$), $t(15)=-4.30$, $p<0.001$, Cohen’s $d=-1.08$.